# KnowEEG: Explainable Knowledge Driven EEG Classification

### v1

@@
Abstract.
Electroencephalography (EEG) is a method of recording brain activity that shows significant promise in applications ranging from disease classification to emotion detection and brain-computer interfaces. Recent advances in deep learning have improved EEG classification performance yet model explainability remains an issue. To address this key limitation of explainability we introduce KnowEEG; a novel explainable machine learning approach for EEG classification. KnowEEG extracts a comprehensive set of per-electrode features, filters them using statistical tests, and integrates between-electrode connectivity statistics. These features are then input to our modified Random Forest model (Fusion Forest) that balances per electrode statistics with between electrode connectivity features in growing the trees of the forest. By incorporating knowledge from both the generalized time-series and EEG-specific domains, KnowEEG achieves performance comparable to or exceeding state-of-the-art deep learning models across five different classification tasks: emotion detection, mental workload classification, eyes open/closed detection, abnormal EEG classification, and event detection. In addition to high performance, KnowEEG provides inherent explainability through feature importance scores for understandable features. We demonstrate by example on the eyes closed/open classification task that this explainability can be used to discover knowledge about the classes. This discovered knowledge for eyes open/closed classification was proven to be correct by current neuroscience literature. Therefore, the impact of KnowEEG will be significant for domains where EEG explainability is critical such as healthcare.
EEG classification, Time Series, Machine Learning, Explainable
ccs: Applied computing Health care information systemsccs: Computing methodologies Machine learning
1. Introduction
@@
High‐level schematic of the two‐thread EEG pipeline. Refer to caption
Figure 1. KnowEEG Pipeline: There are two parallel threads to the pipeline. In thread one 783 features from generalized and EEG specific time series literature are calculated per electrode and concatenated. A two sample Kolmogorov-Smirov test with user specified p-value (default 0.05) is used to filter out uninformative features. In thread two between electrode connectivity features are calculated for several connectivity metrics. The best performing metric is selected using classifier performance on the validation data. The features from thread one and thread two are then combined using the Fusion Forest (see Algorithm 1).
@@
Electroencephalogram (EEG) is a widely used method of recording cortical neuronal activity with a high temporal resolution (casson2018electroencephalogram, ). Data collection is relatively inexpensive and also non-invasive. Thus, EEG has great potential to drive advancements in mental health, enable sophisticated brain-computer interfaces, and enhance disease detection, paving the way for future innovations in healthcare. Much research has already been done on EEG classification with applications ranging from the classification of neurodegenerative diseases such as Alzheimer’s (sahota2024interpretable, ) to motor task recognition for Brain-Computer Interfaces (abiri2019comprehensive, ) and emotion detection (katsigiannis2017dreamer, ).
@@
EEG classification is a challenging task given the low signal-to-noise ratio in EEG signals and the presence of inter-subject variability requiring models to generalize well across individuals. Traditional machine learning methods for EEG classification have relied on standard hand-crafted features as inputs to models such as support vector machines, random forests or neural networks (chaturvedi2017quantitative, ) (betrouni2019electroencephalography, ). These pipelines often use frequency-based features combined with select statistics of the EEG signal (sahota2024interpretable, ). One key limitation for traditional feature-based methods has been poor performance in settings where prior knowledge is limited. For this reason, in the last decade research has focussed more and more on deep learning methodologies.
@@
Deep learning methods in this context range from convolutional neural networks (CNN) applied to the raw EEG data to audio-inspired networks that apply a CNN to EEG spectrograms (time-frequency image of the EEG) (oh2020deep, ). Recent advances in deep learning architectures have also led to transformer-inspired networks such as EEG Conformer (eegconformer, ) and EEG2Rep (mohammadi2024eeg2rep, ) (self-supervised representation-based model) for classification. Although the performance across EEG classification tasks has improved, deep learning models are often opaque and their inner workings are hard to interpret. This leads researchers to post-hoc explainability methods in an attempt to understand models (posthoc_2022, ). Explainability is a particularly critical issue for EEG as it can enhance insights into neurophysiological phenomena, aid clinical decision-making and improve trust in the model.
@@
Therefore, while data- and compute-hungry deep learning methods also have issues around explainability, existing feature-based methods struggle to achieve high classification performance across diverse EEG classification tasks. We revisit the overlooked domain of feature-based methods and present a novel feature based EEG classification pipeline KnowEEG that addresses these limitations achieving both high performance and explainability. We hypothesize that the informative feature space for EEG Classification tasks consists of per-electrode features combined with between-electrode connectivity measures. Thus, we construct KnowEEG to reflect this. We leverage existing time series and EEG knowledge to construct a large feature space of 783 per electrode statistics that are filtered and then combined with between electrode connectivity statistics. These features are input to our modified Random Forest model (Fusion Forest) that balances per electrode statistics with between electrode connectivity features in growing the trees of the forest (see pipeline in Figure 1 and Fusion Forest Algorithm 1). Utilizing a high dimensional feature space enables KnowEEG to achieve high performance across a diverse set of EEG classification tasks. Combining this performance with the explainability of our approach means KnowEEG can also be used to discover new knowledge about the classification task at hand.
@@
2. Related Work
@@
2.1. Feature Engineering for General Time Series
@@
Time series data covers vastly different domains from energy consumption to EEG and weather prediction. Although different, all domains share the fundamental property of sequentially ordered observations that change over time. Basic statistical descriptors such as mean, variance, skewness and kurtosis are often used as a starting point for classification. Features can then be extended to temporal patterns which include autocorrelation at different lags, peak detection and peak counting among others. In (lubba2019catch22, ), the authors propose a set of 22 features for generalized time series classification after extensive testing on 93 time series datasets. The ’catch22’ features include linear and non-linear autocorrelation, successive differences, value distributions and outliers and fluctuation scaling properties (lubba2019catch22, ). Features for generalized time series classification can also include shape-based features such as shapelets or motifs that capture local patterns. Finally, features can also be generated from the frequency domain. These are discussed in detail in 2.2.
2.2. Feature Engineering for EEG
@@
Research has shown frequency-based features to be particularly informative of brain activity (bandpowerpaper, ) and therefore effective for EEG classification. The power spectral density (strength of the EEG signal across frequencies) is calculated using the Fourier transform. From the power spectral density band powers and peak frequency values can be calculated. Band powers refer to specifically defined frequency bands. For our paper we use the band power definitions as delta (0.5–4Hz), theta (4–8Hz), alpha (8–12Hz), sigma (12–16Hz), beta (16–30Hz) and gamma (30–40Hz). Band power values and peak / median frequency features can be input to machine learning models such as support vector machines / random forests for classification as in (chaturvedi2017quantitative, ) and (betrouni2019electroencephalography, ).
@@
The afore-discussed features though effective reduce temporal resolution to the window over which the feature was calculated and discard phase information within the signal. Connectivity measures are metrics for calculating connectivity between regions of the brain and often utilize phase information. Phase Lag Index (PLI) for example (kuang2022phase, ), measures the asymmetry of the distribution of phase differences across time between two signals with its value ranging between 0 and 1. A PLI of zero means no coupling (or instantaneous coupling) and a value of 1 means true lagged interaction. Connectivity measures can be calculated for the entire EEG signal between electrodes or for sub-components of the EEG signal within the frequency domain (commonly band powers). For example in (kuang2022phase, ), researchers find the phase lag index in the alpha band to be correlated with cognitive assessment score in the task of Mild Cognitive Impairment classification.
@@
Brain connectivity in general comprises the following subcomponents; structural connectivity (anatomical connectivity between brain regions), functional connectivity (statistical dependencies or correlations in activity) and effective connectivity (causal or directional influences between regions) (chiarion2023connectivity, ). Another commonly used metric is coherence which measures functional connectivity via correlation in the frequency domain. In (waninger2020neurophysiological, ) researchers use coherence combined with other signal statistics to achieve over 90% accuracy in the task of Parkinson’s classification.
@@
Thus, in summary, feature-based approaches inspired by domain knowledge have been successful for EEG classification often focusing on frequency-based properties of the signal and measures of connectivity across the brain. However, approaches vary from task to task and there is yet to be the development of a generalized EEG feature engineering-based classification pipeline across task domains.
2.3. Deep Learning for EEG
@@
Over the last decade, deep learning has experienced unprecedented growth and success across domains. This has led to much research experimenting with existing deep learning architectures applied to EEG data and the development of EEG-specific deep neural architectures(mohammadi2024eeg2rep, ).
@@
Firstly, researchers have effectively applied simple convolutional neural networks (CNNs) directly to processed EEG data for classification. In (oh2020deep, ) a CNN comprising of a series of 1D-convolutions and max pooling layers achieved just under 90% accuracy on Parkinson’s classification. EEG researchers have since taken inspiration from the audio domain by applying CNNs to time-frequency images (spectrograms) of the raw EEG signal (li2022deep, ) (mandhouj2021automated, ).
@@
Deep learning architectures have also been developed to deal directly with sequential data from RNN variants such as LSTMs and GRU’s to more recently Transformers. Transformers excel at capturing long-range dependencies and as such have been successfully used for EEG classification. In (song2022eeg, ) the authors present a novel architecture that uses a convolutional module to generate embeddings for input EEG data which is then passed to a Transformer Encoder module for classification. This EEG Conformer achieves state of the art performance on many brain computer interface based tasks (song2022eeg, ).
@@
One issue with deep neural networks is susceptibility to over-fitting (mohammadi2024eeg2rep, ). With EEG data this is particularly problematic as signals are noisy and inter-subject variability can be present (schirrmeister2017deep, ). Self-supervised learning (SSL) aims to solve this problem by learning a representation of the EEG and generating a self-supervisory signal from the data (weng2024self, ). These models can learn from labelled or unlabeled data and the learnt representations used for classification. State-of-the-art SSL deep learning models include BENDR (kostas2021bendr, ), BIOT (yang2024biot, ) and most recently EEG2Rep (mohammadi2024eeg2rep, ).
@@
3. Methodology of KnowEEG
3.1. Problem Statement
@@
3.2. KnowEEG Model Architecture
@@
Per electrode we calculate 783 features from both the generalized time series domain and the EEG-specific domain. This includes basic features like mean, standard deviation and kurtosis to more complex time series features like autocorrelation at different lags and EEG specific features related to frequency components present in the signal. In order to do this calculation the time series Python package TSFresh (christ2018time, ) is used. Exact settings for feature calculation are presented in 4.3 and further details for implementation provided in A. These per electrode features are concatenated after being calculated. Therefore, for our 14 electrode datasets there are 783 features x 14 electrodes which equals 10,962 time series features per sample (12,528 for 16 electrode datasets). Our aim is to create a broad set of time series descriptors for each EEG sample. We hypothesize that this broad set of features contains the subset of informative features for the specific EEG classification task. To avoid model over-fitting and reduce unnecessary computation we then filter these features using the Two-sample Kolmogorov–Smirnov (KS) test (for binary classification) or Kruskal–Wallis hypothesis test (for multi-class). The p-value of 0.05 is used as default for these tests. Thus, the per electrode thread of the pipeline results in a final set of filtered per electrode features as shown in Figure 1.
@@
We calculate a separate set of between electrode connectivity features. We select Spearman and Pearson correlation, which measure linear and non-linear correlation between electrodes. For a 14 electrode dataset this results in 182 correlation features (91 Spearman and 91 Pearson). We also select a series of well researched EEG specific connectivity measures. For our pipeline we select Coherence (Coh) , Imaginary Coherence (ImCoh), Pairwise Phase Consistency (PPC), Phase Lag Value (PLV), Phase Lag Index (PLI) , Directed Phase Lag Index (DPLI) and Weighted Phase Lag Index (WPLI) (MNE_2013, ). Users of the pipeline can select any number of connectivity metrics that they expect could be informative for classification. These connectivity metrics can be calculated for the entire signal or on sub components of the signal. We propose calculating these metrics for each sub band of the EEG signal with regards to the bands defined in 2.2. Therefore per connectivity metric, for the 14 electrode datasets there are 91 (features per sub band) x 6 sub bands (delta, theta, alpha, sigma, beta, gamma) which equates to 546 total features per metric (720 features for 16 electrode datasets). Some connectivity metrics require segmentation of the raw signal into epochs prior to calculation. We follow standard protocol for segmentation and detail this in Appendix B. Therefore we have 8 candidate connectivity metrics (correlation - Spearman and Pearson) aswell as the afore stated 7 specific EEG metrics.
@@
We now have per electrode features from thread one and between electrode features (connectivity measure) from thread two of the pipeline. We consider this data to be of two different ’modes’ and thus require a method of combining the information from both modes. For a classifier we select the Random Forest as it is inherently explainable and has been shown to perform well on EEG classification from features (edla2018classification, ) (fraiwan2012automated, ). For fusion of the two modes we can select feature level fusion, decision level fusion or propose another method. As the number of features per electrode can be far higher than the number of connectivity metrics we decide against feature level fusion as for tree-based models (especially Random Forests) this can skew the model towards the mode with more features. Decision-level fusion would be viable. However, we hypothesize that highest performance will be achieved if each tree in the Random Forest has access to features from both modes of data. We thus propose a modification of the traditional Random Forest algorithm so that for each tree a random subset of features is selected from mode 1 and separately from mode 2. We have not seen this simple modification to the Random Forest for feature fusion in the literature and therefore we outline the algorithm for this in 1. We call this the Fusion Forest.
Algorithm 1 Fusion Forest
1:  Input: $N$ samples, $Xa$ with $A$ total features (mode 1), $Xb$ with $B$ total features (mode 2)
@@
4:     Select random subset of samples $N_{\text{subset}}$ (bootstrap sample of size $N$)
5:     Select a random subset of $\sqrt{A}$ features from $Xa$ giving $Xa_{\text{subset}}$
6:     Select a random subset of $\sqrt{B}$ features from $Xb$ giving $Xb_{\text{subset}}$
@@
4. Experiments and Discussion
4.1. Datasets
We use five publicly available EEG datasets. Three of the datasets recorded data with 14 channel Emotiv EEG headsets. These datasets are DREAMER (emotion detection) (katsigiannis2017dreamer, ) , STEW (Mental workload Classification) (lim2018stew, ) and Crowdsourced (eyes open / close detection) (williams2023crowdsourced, ). EEG from these datasets had a sampling rate of 128Hz upon recording and all samples are two seconds long (after pre-processing). Recorded data for these three datasets is processed as per (mohammadi2024eeg2rep, ), snipping the data to create 2 second segments. Further details are provided in C.
The other two datasets are from the Temple University Hospital (TUH) Corpus (TUAB_dataset, ) (TUEV_dataset, ). TUH is one of the largest EEG data repositories in the world with data collected in a lab setting with a range of different EEG recording devices. In processing this data we select 16 standard EEG channels following the 10-20 international system. This data was recorded at 256Hz with 5 second samples for TUEV (TUEV_dataset, ) and 10 second samples for TUAB (TUAB_dataset, ). We provide further details of processing in Appendix C.
The Emotiv datasets were split into train / validation / test sets subject wise, challenging models to learn generalizable patterns across subjects. The TUH datasets (TUEV and TUAB) were inherently split already into train and test sets. We further split the train sets into 80% training and 20% validation. Again, this is following the same protocol as in (mohammadi2024eeg2rep, ).
Table 1 shows the main characteristics of all five data sets. Further information on each dataset and pre-processing can be found in Appendix C.
Table 1. Properties of the EEG datasets used (adapted from (mohammadi2024eeg2rep, )). DREAMER (katsigiannis2017dreamer, ), STEW (lim2018stew, ), Crowdsourced (williams2023crowdsourced, ), TUEV (TUEV_dataset, ) and TUAB (TUAB_dataset, )

### v2

@@
Abstract
Electroencephalography (EEG) is a method of recording brain activity that shows significant promise in applications ranging from disease classification to emotion detection and brain-computer interfaces. Recent advances in deep learning have improved EEG classification performance yet model explainability remains limited. To address this key limitation we introduce KnowEEG; a novel explainable machine learning approach for EEG classification. KnowEEG extracts a comprehensive set of per-electrode features, filters them using statistical tests, and integrates between-electrode connectivity statistics. These features are then input to our modified Random Forest model (Fusion Forest) that balances per electrode statistics with between electrode connectivity features in growing the trees of the forest. By incorporating knowledge from both the generalized time-series and EEG-specific domains, KnowEEG achieves performance comparable to or exceeding state-of-the-art deep learning models across five different classification tasks: emotion detection, mental workload classification, eyes open/closed detection, abnormal EEG classification, and event detection. In addition to high performance, KnowEEG provides inherent explainability through feature importance scores for understandable features. We demonstrate by example on the eyes closed/open classification task that this explainability can be used to discover knowledge about the classes. This discovered knowledge for eyes open/closed classification was proven to be correct by current neuroscience literature. Therefore, the impact of KnowEEG will be significant for domains where EEG explainability is critical such as healthcare.
journal: Neurocomputing
@@ Abstract
[inst1]organization=University of Bristol, city=Bristol, country=United Kingdom
[…]
[inst2]organization=Monash University, city=Melbourne, state=Victoria, country=Australia
1 Introduction
@@
Electroencephalogram (EEG) is a widely used method of recording cortical neuronal activity with a high temporal resolution [7]. Data collection is relatively inexpensive and also non-invasive. Thus, EEG has great potential to drive advancements in mental health, enable sophisticated brain-computer interfaces, and enhance disease detection, paving the way for future innovations in healthcare. Extensive research has explored EEG classification with applications ranging from the classification of neurodegenerative diseases such as Alzheimer’s [35] to motor task recognition for Brain-Computer Interfaces [2] and emotion detection [24].
@@
EEG classification is a challenging task given the low signal-to-noise ratio in EEG signals and the presence of inter-subject variability requiring models to generalize well across individuals. Traditional machine learning methods for EEG classification have relied on standard hand-crafted features as inputs to models such as support vector machines, random forests or neural networks [8] [6]. These pipelines often use frequency-based features combined with select statistics of the EEG signal [35]. One key limitation for traditional feature-based methods has been poor performance in settings where prior knowledge is limited. For this reason, in the last decade research has focussed increasingly on deep learning methodologies.
@@
Deep learning methods in this context range from convolutional neural networks (CNN) applied to the raw EEG data to audio-inspired networks that apply a CNN to EEG spectrograms (time-frequency image of the EEG) [33]. Recent advances in deep learning architectures have also led to transformer-inspired networks such as EEG Conformer [38] and EEG2Rep [32] (self-supervised representation-based model) for classification. Although the performance across EEG classification tasks has improved, deep learning models are often opaque and their inner workings are hard to interpret. This leads researchers to post-hoc explainability methods in an attempt to understand models [22]. Explainability is a particularly critical issue for EEG as it can enhance insights into neurophysiological phenomena, aid clinical decision-making and improve trust in the model.
@@
Therefore, while data and compute hungry deep learning methods also have issues around explainability, existing feature-based methods struggle to achieve high classification performance across diverse EEG classification tasks. We revisit the overlooked domain of feature-based methods and present a novel feature based EEG classification pipeline KnowEEG that addresses these limitations achieving both high performance and explainability. We hypothesize that the informative feature space for EEG Classification tasks consists of per-electrode features combined with between-electrode connectivity measures. Thus, we construct KnowEEG to reflect this. We leverage existing time series and EEG knowledge to construct a large feature space of 783 per electrode statistics that are filtered and then combined with between electrode connectivity statistics. These features are input to our modified Random Forest model (Fusion Forest) that balances per electrode statistics with between electrode connectivity features in growing the trees of the forest (see pipeline in Figure 1 and Fusion Forest Algorithm 1). Utilizing a high dimensional feature space enables KnowEEG to achieve high performance across a diverse set of EEG classification tasks. Combining this performance with the explainability of our approach means KnowEEG can also be used to discover new knowledge about the classification task at hand.
@@
Figure 1: KnowEEG Pipeline: There are two parallel threads to the pipeline. In thread one, 783 features from generalised and EEG-specific time-series literature are calculated per electrode and concatenated. Feature relevance is evaluated via univariate non-parametric hypothesis tests (Mann-Whitney U for binary targets and Kruskal-Wallis tests for multiclass targets), with Benjamini–Yekutieli correction to control the false discovery rate ($\alpha=0.05$), filtering out uninformative features. In thread two, between-electrode connectivity features are calculated for connectivity metrics. The best-performing metric is selected using classifier performance on the validation data. The features from thread one and thread two are then combined using the Fusion Forest (see Algorithm 1).
2 Related Work
@@
2.1 Feature Engineering for General Time Series
@@
Time series data covers vastly different domains from energy consumption to EEG and weather prediction. Although different, all domains share the fundamental property of sequentially ordered observations that change over time. Basic statistical descriptors such as mean, variance, skewness and kurtosis are often used as a starting point for classification. Features can then be extended to temporal patterns which include autocorrelation at different lags, peak detection and peak counting among others. In [30], the authors propose a set of 22 features for generalized time series classification after extensive testing on 93 time series datasets. The ‘catch22’ features include linear and non-linear autocorrelation, successive differences, value distributions and outliers and fluctuation scaling properties [30]. Features for generalized time series classification can also include shape-based features such as shapelets or motifs that capture local patterns. Finally, features can also be generated from the frequency domain. These are discussed in detail in 2.2.
2.2 Feature Engineering for EEG
@@
Research has shown frequency-based features to be particularly informative of brain activity [20] and therefore effective for EEG classification. The power spectral density (strength of the EEG signal across frequencies) is calculated using the Fourier transform. From the power spectral density band powers and peak frequency values can be calculated. Band powers refer to specifically defined frequency bands. For our paper we use the band power definitions as delta (0.5–4Hz), theta (4–8Hz), alpha (8–12Hz), sigma (12–16Hz), beta (16–30Hz) and gamma (30–40Hz). Band power values and peak / median frequency features can be input to machine learning models such as support vector machines or random forests for classification as in [8] and [6].
@@
The afore-discussed features though effective reduce temporal resolution to the window over which the feature was calculated and discard phase information within the signal. Connectivity measures are metrics for calculating connectivity between regions of the brain and often utilize phase information. Phase Lag Index (PLI) for example [26], measures the asymmetry of the distribution of phase differences across time between two signals with its value ranging between 0 and 1. A PLI of zero means no coupling (or instantaneous coupling) and a value of 1 means true lagged interaction. Connectivity measures can be calculated for the entire EEG signal between electrodes or for sub-components of the EEG signal within the frequency domain (commonly band powers). For example in [26], researchers find the phase lag index in the alpha band to be correlated with cognitive assessment score in the task of Mild Cognitive Impairment classification.
@@
Brain connectivity in general comprises the following subcomponents; structural connectivity (anatomical connectivity between brain regions), functional connectivity (statistical dependencies or correlations in activity) and effective connectivity (causal or directional influences between regions) [9]. Another commonly used metric is coherence which measures functional connectivity via correlation in the frequency domain. In [39] researchers use coherence combined with other signal statistics to achieve over 90% accuracy in the task of Parkinson’s classification.
@@
Thus, in summary, feature-based approaches inspired by domain knowledge have been successful for EEG classification often focusing on frequency-based properties of the signal and measures of connectivity across the brain. While many feature-engineering approaches exist, none provide a unified, high-performance, explainable pipeline applicable across multiple EEG classification domains.
2.3 Deep Learning for EEG
@@
Over the last decade, deep learning has experienced unprecedented growth and success across domains. This has led to much research experimenting with existing deep learning architectures applied to EEG data and the development of EEG-specific deep neural architectures[32].
@@
Firstly, researchers have effectively applied simple convolutional neural networks (CNNs) directly to processed EEG data for classification. In [33] a CNN comprising of a series of 1D-convolutions and max pooling layers achieved just under 90% accuracy on Parkinson’s classification. EEG researchers have since taken inspiration from the audio domain by applying CNNs to time-frequency images (spectrograms) of the raw EEG signal [27] [31].
@@
Deep learning architectures have also been developed to deal directly with sequential data from RNN variants such as LSTMs and GRU’s to more recently Transformers. Transformers excel at capturing long-range dependencies and as such have been successfully used for EEG classification. In [37] the authors present a novel architecture that uses a convolutional module to generate embeddings for input EEG data which is then passed to a Transformer Encoder module for classification. This EEG Conformer achieves state of the art performance on many brain computer interface based tasks [37].
@@
One issue with deep neural networks is susceptibility to over-fitting [32]. With EEG data this is particularly problematic as signals are noisy and inter-subject variability can be present [36]. Self-supervised learning (SSL) aims to solve this problem by learning a representation of the EEG and generating a self-supervisory signal from the data [40]. These models can learn from labelled or unlabeled data and the learnt representations used for classification. State-of-the-art SSL deep learning models include BENDR [25], BIOT [42] and most recently EEG2Rep [32].
@@
3 Methodology of KnowEEG
3.1 Problem Statement
@@
3.2 KnowEEG Model Architecture
@@
Per electrode we calculate 783 features from both the generalized time series domain and the EEG-specific domain. This includes basic features like mean, standard deviation and kurtosis to more complex time series features like autocorrelation at different lags and EEG specific features related to frequency components present in the signal. In order to do this calculation the time series Python package TSFresh [12] is used. Exact settings for feature calculation are presented in 4.3 and further details for implementation provided in A. These per electrode features are concatenated after being calculated. Therefore, for our 14 electrode datasets there are 783 features x 14 electrodes which equals 10,962 time series features per sample (12,528 for 16 electrode datasets). Our aim is to create a broad set of time series descriptors for each EEG sample. We hypothesize that this broad set of features contains the subset of informative features for the specific EEG classification task. To avoid model over-fitting and reduce unnecessary computation, features are filtered using univariate non-parametric hypothesis tests to assess their relevance to the classification target. For binary classification we use the Mann-Whitney U test while for multi-class classification Kruskal–Wallis tests are applied. Resulting p-values are adjusted to control the false discovery rate using the Benjamini-ekutieli correction, with a significance threshold of 0.05 [5]. Thus, the per electrode thread of the pipeline results in a final set of filtered per electrode features as shown in Figure 1.
@@
We calculate a separate set of between electrode connectivity features. We select Spearman and Pearson correlation, which measure linear and non-linear correlation between electrodes. For a 14 electrode dataset this results in 182 correlation features (91 Spearman and 91 Pearson). We also select a series of well researched EEG specific connectivity measures. For our pipeline we select Coherence (Coh) , Imaginary Coherence (ImCoh), Pairwise Phase Consistency (PPC), Phase Lag Value (PLV), Phase Lag Index (PLI) , Directed Phase Lag Index (DPLI) and Weighted Phase Lag Index (WPLI) [18]. Users of the pipeline can select any number of connectivity metrics that they expect could be informative for classification. These connectivity metrics can be calculated for the entire signal or on sub components of the signal. We propose calculating these metrics for each sub band of the EEG signal with regards to the bands defined in 2.2. Therefore per connectivity metric, for the 14 electrode datasets there are 91 (features per sub band) x 6 sub bands (delta, theta, alpha, sigma, beta, gamma) which equates to 546 total features per metric (720 features for 16 electrode datasets). These connectivity metrics require segmentation of the raw signal into epochs prior to calculation. We follow standard protocol for segmentation and detail this in C. In addition to these conventional connectivity metrics we also define Functional Power Connectivity (FPC) which consists of per electrode relative band powers and electrode-electrode features capturing spatial differences in oscillatory power across brain regions. Unlike phase-based metrics, FPC can be computed reliably even from short signals where windowing introduces noise. See B for full FPC definition. Therefore, in total we have 9 candidate connectivity metrics.
@@
We now have per electrode features from thread one and between electrode features (connectivity measure) from thread two of the pipeline. We consider this data to be of two different ‘modes’ and thus require a method of combining the information from both modes. For a classifier we select the Random Forest as it is inherently explainable and has been shown to perform well on EEG classification from features [14] [15]. For fusion of the two modes we can select feature level fusion, decision level fusion or propose another method. As the number of features per electrode can be far higher than the number of connectivity metrics we decide against feature level fusion as for tree-based models (especially Random Forests) this can skew the model towards the mode with more features. Decision-level fusion would be viable. However, we hypothesize that highest performance will be achieved if each tree in the Random Forest has access to features from both modes of data. We thus propose a modification of the traditional Random Forest algorithm so that for each tree a random subset of features is selected from mode 1 and separately from mode 2. We have not seen this simple modification to the Random Forest for feature fusion in the literature and therefore we outline the algorithm for this in 1. We call this the Fusion Forest.
Algorithm 1 The Fusion Forest Algorithm. This modified Random Forest balances two feature modes ($X_{a}$ and $X_{b}$) by independently selecting a random subset of features from each (lines 5-6) before fitting each of the $K$ trees. This prevents the feature set with a higher dimension from dominating the model.
1:Input: NN samples, XaXa with AA total features (mode 1), XbXb with $B$ total features (mode 2)
@@
4:  Select random subset of samples $N_{\text{subset}}$ (bootstrap sample of size NN)
5:  Select a random subset of $\sqrt{A}$ features from XaXa giving $Xa_{\text{subset}}$
6:  Select a random subset of $\sqrt{B}$ features from XbXb giving $Xb_{\text{subset}}$
@@
4 Experiments and Discussion
4.1 Datasets
We use five publicly available EEG datasets. Three of the datasets recorded data with 14 channel Emotiv EEG headsets. These datasets are DREAMER (emotion detection) [24] , STEW (Mental workload Classification) [28] and Crowdsourced (eyes open / close detection) [41]. EEG from these datasets had a sampling rate of 128Hz upon recording and all samples are two seconds long (after pre-processing). Recorded data for these three datasets is processed as per [32], snipping the data to create 2 second segments. Further details are provided in D.
The other two datasets are from the Temple University Hospital (TUH) Corpus [29] [21]. TUH is one of the largest EEG data repositories in the world with data collected in a lab setting with a range of different EEG recording devices. In processing this data we select 16 standard EEG channels following the 10-20 international system. This data was recorded at 256Hz with 5 second samples for TUEV [21] and 10 second samples for TUAB [29]. We provide further details of processing in Appendix D.
The Emotiv datasets were split into train / validation / test sets subject wise, challenging models to learn generalizable patterns across subjects. The TUH datasets (TUEV and TUAB) were inherently split already into train and test sets. We further split the train sets into 80% training and 20% validation. Again, this is following the same protocol as in [32].
Table 1 shows the main characteristics of all five data sets. Further information on each dataset and pre-processing can be found in D.
Table 1: Properties of the EEG datasets used (adapted from [32]). DREAMER [24], STEW [28], Crowdsourced [41], TUEV [21] and TUAB [29]